# Chain-of-Authorization: Internalizing Authorization into Large Language Models via Reasoning Trajectories

### 24 Mar 2026 v1

@@ 5.2 Methodology
Output reasoning trajectory reconstruction: After input reconstruction, the model needs to generate an explicit authorization inference trajectory $T$ as a causal prefix generated by the response $O$. The authorization chain $T$ consists of three highly coupled stages:
$T=(T_{res},T_{id},T_{dec})$ (9)
where $T_{res}$ represents the resource review stage, the LLM first parses the user’s intent to query $X$ and infers the permission label $C_{x}$ of the endogenous or exogenous knowledge required to answer the query. It also parses the context label $C_{e}$ to clarify the permissions required for the corresponding context. $T_{id}$ denotes the identity resolution stage, in which the LLM extracts the user’s permission $C_{u}$ from the system prompt. If tool calls are involved, the model also needs to parse the specific tool permission $C_{g}$, thereby establishing the access boundaries of the permission subject. Finally, $T_{dec}$ represents the decision-making process. Based on the parsing results of $T_{res}$ and $T_{id}$, the model applies the policy function to derive the authorization decision for the user’s current query. This structured sequence-enforcing model must first establish a logically consistent compliance proof before generating the final response.
Training data synthesis and paradigm alignment: To internalize the aforementioned reasoning capabilities within the model, we designed an authorization-aware structured training paradigm.
First, based on the extent to which user permissions cover the permissions required for the model to generate compliant responses, we constructed three types of sample sequences $(C,X,T,Y)$ that cover different authorization states by permuting user permissions and model responses. This allows the model to learn fine-grained permission matching relationships:
  • Matched authorization: In the matched authorization state, $C_{u}$ covers all permission required for responding prompt $X$. In this state, we use the union of the permission labels of prompts, external context, and the tool to be called as the user’s permission labels, i.e., $\{C_{e}\cup C_{x}\cup C_{g}\}=C_{u}$. The model responds normally to the user prompt.
Secondly, we adopt a unified supervised fine-tuning (SFT) framework. By minimizing the negative log-likelihood loss under the same sequence prediction objective, gradient descent is used to bind authorization decisions and response generation:
$\mathcal{L}(\theta)=-\sum_{(C,X,T,Y)\in\mathcal{D}}\log\pi_{\theta}(Y,T|X,E,C)$ (10)
It is worth mentioning that we did not design complex multi-task objectives or auxiliary loss functions. This general loss function makes CoA a completely data-driven approach, independent of specific model architectures, and therefore widely applicable to various mainstream LLMs.
@@ 5.3 Experiment Setup and Evaluation
5.3.1 Multidimensional Benchmarking Scenarios
We divided the evaluation task into three scenarios: internal parameterized knowledge, external context, and tool invocation, to verify the generality of CoA in handling heterogeneous information flows. In the internal parameterized knowledge control scenario, we used the WMDP [li2024wmdp] and MMLU [hendryckstest2021mmlu] datasets. In the external context knowledge control scenario, we used the SQuAD and COVID-QA datasets. Finally, in the tool calling control scenario, we used the Mobile-Actions dataset. Information about the datasets used is described below:
@@ 5.3.1 Multidimensional Benchmarking Scenarios
  • Mobile-Actions [google_mobile_actions]: A mobile operation instruction understanding and execution dataset released by Google. We construct seven access labels across two dimensions: object type (system, information, communication) and operation permissions (read/write). We assign two access labels to each tool, corresponding to the two dimensions mentioned above. This design aims to simulate the limitations of tool access under different authorization states.
Detailed dataset statistics are provided in Table 4.
Table 4: Dataset Details.
WMDP MMLU SQuAD Covid-QA Mobile-Actions
#Permission Labels 3 57 100 1234 5
@@ 5.3.2 Backbone Models and Baseline methods
  • PermissionLM [jayaraman2025permlm]: A structured isolation method that blocks illegal information flow through physical sharding or training independent sub-models.
  • SudoLM [liu2025sudolm]: A prompt-guided method that guides the model to execute access policies solely through permission tags embedded in system prompts.
  • External Gateway: Simulates a two-phase audit architecture, using an external model to first determine permissions and then decide whether to allow the main model to generate a response. In our experiments, we fine-tuned Roberta-base [roberta] on the aforementioned datasets, using prompts as input and permission labels as the target.
To evaluate CoA’s robustness to adversarial prompts, we also conducted robustness experiments on the WMDP dataset. This dataset contains sensitive knowledge related to biological, cyber, and chemical security, making it an ideal scenario for testing the model’s authorization boundaries. We used three representative adversarial attack methods to induce the model to output unauthorized content:
@@
5.3.3 Evaluation Metrics and Implementation Details
@@
  • Security measures the model’s ability to correctly trigger a rejection response in unauthorized scenarios. For statistical convenience, we require all rejected responses sampled in $\delta_{rej}$ to begin with ”Sorry”. We determine whether the model rejected the question by counting the cases where the final response contains ”Sorry”. At last, we use the rejection rate, the percentage of prompts for which the model rejected the question out of all prompts, to reflect security.
Furthermore, in robustness experiments, we used the attack success rate to measure the frequency with which attack methods induce controlled knowledge leakage from the model. Since we require all rejected responses sampled in $\delta_{rej}$ to begin with ”Sorry”, we determined whether the attack method bypassed permission checks by counting cases where the final response did not contain ”Sorry”. Finally, we used the percentage of prompts that bypassed permission checks out of all prompts to reflect the attack success rate.
Implementation Details. All methods were implemented within a unified, parameter-efficient fine-tuning paradigm. We used Low-Rank Adaptation (LoRA) [hu2022lora] to fine-tune the backbone model, applying the LoRA module to all linear layers and uniformly setting the rank to $64$.
All models are trained iteratively for 3 epochs on the full training set. To support the newly added permission labels, we expand the vocabulary and embedding layer size of the backbone model during the training of SudoLM and CoA. This improvement ensures that permission labels have an independent semantic representation space, avoiding conflicts with the representation of general vocabulary. We searched for the optimal learning rate for different methods and datasets during training. For SFT, PermissionLM, and CoA, we use a learning rate of $1.0\texttimes 10^{-4}$ on the WMDP, MMLU, and Mobile-Actions datasets; and adjust it to $5.0\texttimes 10^{-4}$ on SQuAD and COVID-QA, which involve long contexts. For SudoLM, to maintain a balanced prompt guidance, we uniformly use a more conservative learning rate of $5.0\texttimes 10^{-6}$ across all datasets.
@@
6 Related Work
As LLMs become the cognitive core of AI systems like autonomous agents, effectively constraining their ability to access internal knowledge and external tools has become a key safety challenge. Existing methods to achieve this goal primarily fall into three categories: structural isolation, prompt guidance, and safety alignment.
Structural isolation methods attempt to limit the information the model can access through physical boundaries. For example, Tiwari et al. [tiwari2024ifc] and Jayaraman et al. [jayaraman2025permlm] proposed training separate submodels for data with different access levels and using a modular structure to support users with multiple permissions. DOMBA [segal2025domba] prevents unauthorized access through structured isolation and aggregation of independent model outputs. In the Retrieval Augmentation (RAG) scenario, Wutschitz et. al. [wutschitz2023rethinking] proposed directly blocking unauthorized documents during the retrieval phase of retrieval enhancement generation. However, this paradigm suffers from severe scalability bottlenecks. When dealing with fine-grained, dynamically changing permission relationships in enterprise environments, physical isolation not only incurs extremely high computational and storage costs but also requires model retraining whenever permission rules change. In contrast, CoA internalizes access control into a single model’s inference capability, enabling highly scalable dynamic authorization without introducing additional models.
Prompt guidance methods aim to constrain model generation by explicitly introducing conditional information at the input. Early conditional training methods (such as CTRL [keskar2019ctrl], Prefix-tuning [li2021prefix], and PPLM [dathathri2020pplm]) primarily focused on controlling the output format and style, without considering strict access constraints. Recent prompt-guided methods, such as SudoLM [liu2025sudolm], SudoLLM [saha2025sudollm], and role-aware LLMs [almheiri2025role], are attempting to embed user permission credentials into system cue words to impose soft constraints on large models. While these methods avoid the overhead of repeatedly training multiple models, they are prone to cognitive confusion in complex scenarios such as valid yet incorrect authorization status [almheiri2025role], making them easily bypassed by adversarial inputs. Our proposed CoA, by explicitly introducing authorization inference trajectories, transforms authorization into a causal premise for the generation process, preventing a disconnect between policy and inference and thereby allowing LLMs to discriminate fine-grained authorization status.
Furthermore, some methods rely on privacy-preserving techniques and safety alignment to limit the output of sensitive information by LLMs. For example, differential privacy fine-tuning [dp-ft] and piecewise aggregation training reduce the model’s memory of sensitive data by limiting the gradient of individual samples; machine forgetting techniques modify weights to completely erase specific privacy information from the model; and alignment paradigms such as reinforcement learning based on human feedback (RLHF) use reward signals to encourage the model to learn to refuse sensitive requests. However, these methods tend to globally suppress sensitive information and are difficult to support identity-based differentiated access. In contrast, CoA does not rely on knowledge unlearning but learns permission boundaries during inference, achieving fine-grained access control that is ”visible on demand”.
7 Data availability
All models and datasets used in our experiments are open-sourced and are available on Huggingface444huggingface.com.
8 Code availability

### 03 Apr 2026 v2

@@ 5.1 Permission-aware Data Synthesis
We designed specialized authorization trajectory templates for the various scenarios tested in our experiments. An extremely simple illustration of an internal parametric knowledge authorization scenario is shown in Figure 6, where the $2$ to $4$ lines represent the resource review, identity resolution, and decision-making stages, respectively. Illustrations of external context authorization and tool-calling authorization scenarios are shown in Figure 7. During synthesis, we generate trajectories for different authorization states by substituting permission labels in the [User Permission] field and updating the corresponding [Final Decision] field.
@@ 5.2 Experiment Setup and Evaluation
5.2.1 Multidimensional Benchmarks
Table 3: Dataset Details.
WMDP MMLU SQuAD Covid-QA Mobile-Actions
#Permission Labels 3 57 100 1234 5
We divided the evaluation task into three scenarios: internal parameterized knowledge, external context, and tool invocation, to verify the generality of CoA in handling heterogeneous information flows. Detailed dataset statistics are provided in Table Table 3. In the internal parameterized knowledge control scenario, we used the WMDP [li2024wmdp] and MMLU [hendryckstest2021mmlu] datasets. In the external context knowledge control scenario, we used the SQuAD and COVID-QA datasets. Finally, in the tool calling control scenario, we used the Mobile-Actions dataset. Information about the datasets used is described below:
@@ 5.2.1 Multidimensional Benchmarks
  • Mobile-Actions [google_mobile_actions]: A mobile operation instruction understanding and execution dataset released by Google. We construct seven access labels across two dimensions: object type (system, information, communication) and operation permission (read/write). We assign two permission labels to each tool, corresponding to the two dimensions mentioned above. This design aims to simulate the limitations of tool access under different authorization states.
@@ 5.2.2 Backbone Models and Baseline methods
  • Permissioned LLM [jayaraman2025permlm]: A structured isolation method that blocks illegal information flow through physical sharding or training independent sub-models.
  • SudoLM [liu2025sudolm]: A prompt-guided method that guides the model to execute access policies solely through permission labels embedded in system prompts.
  • External Gateway: Simulates a two-phase audit architecture, using an external model to first determine permissions and then decide whether to allow the main model to generate a response. In our experiments, we fine-tuned the Roberta-base [roberta] on the aforementioned datasets, using prompts as input and permission labels as the target.
To evaluate CoA’s robustness against adversarial prompts, we also conducted robustness experiments on the WMDP dataset. This dataset contains sensitive knowledge related to biological, cyber, and chemical security, making it an ideal scenario for testing the model’s authorization boundaries. We used three representative adversarial attack methods to induce the model to output unauthorized content:
@@
5.2.3 Evaluation Metrics
@@
  • Security measures the model’s ability to correctly trigger a rejection response in unauthorized scenarios. For statistical convenience, we require all rejected responses sampled in $\delta_{rej}$ to begin with “Sorry”. We determine whether the model rejected the question by counting the cases where the final response contains “Sorry”. At last, we use the rejection rate, the percentage of prompts for which the model rejected the question out of all prompts, to reflect security.
Furthermore, in robustness experiments, we used the attack success rate to measure the frequency with which attack methods induce controlled knowledge leakage from the model. Since we require all responses sampled in $\delta_{rej}$ to begin with “Sorry”, we determined whether the attack method bypassed permission checks by counting cases in which the final response did not contain “Sorry”. Finally, we used the percentage of prompts that bypassed permission checks out of all prompts to reflect the attack success rate.
5.2.4 Implementation Details
All methods were implemented within a unified, parameter-efficient fine-tuning paradigm. We used Low-Rank Adaptation (LoRA) [hu2022lora] to fine-tune the backbone model, applying the LoRA module to all linear layers and setting the rank to $64$ uniformly across all layers.
All models are trained iteratively for $3$ epochs on the full training set. To support the newly added permission labels, we expand the vocabulary and embedding layer size of the backbone model during the training of SudoLM and CoA. This improvement ensures that permission labels have an independent semantic representation space, avoiding conflicts with the representation of general vocabulary. We searched for the optimal learning rate for different methods and datasets during training. For SFT, PermLM, and CoA, we use a learning rate of $1.0\times 10^{-4}$ on the WMDP, MMLU, and Mobile-Actions datasets; and adjust it to $5.0\times 10^{-4}$ on SQuAD and COVID-QA, which involve long contexts. For SudoLM, we uniformly use a more conservative learning rate of $5.0\times 10^{-6}$ on all datasets.
For SoduLM and CoA, we add permission labels to the model’s vocabulary as new tokens to prevent token fragmentation.